### v1

# Jason Zhaoxin Fan3, Bo Tang2, Shichao Song1, Mengwei Wang1, Jiawei Yang1

Abstract
The indexing-retrieval-generation paradigm of retrieval-augmented generation (RAG) has been highly successful in solving knowledge-intensive tasks by integrating external knowledge into large language models (LLMs). However, the incorporation of external and unverified knowledge increases the vulnerability of LLMs because attackers can perform attack tasks by manipulating knowledge. In this paper, we introduce a benchmark named SafeRAG designed to evaluate the RAG security. First, we classify attack tasks into silver noise, inter-context conflict, soft ad, and white Denial-of-Service. Next, we construct RAG security evaluation dataset (i.e., SafeRAG dataset) primarily manually for each task. We then utilize the SafeRAG dataset to simulate various attack scenarios that RAG may encounter. Experiments conducted on 14 representative RAG components demonstrate that RAG exhibits significant vulnerability to all attack tasks and even the most apparent attack task can easily bypass existing retrievers, filters, or advanced LLMs, resulting in the degradation of RAG service quality. Code is available at: https://github.com/IAAR-Shanghai/SafeRAG.
SafeRAG: Benchmarking Security in Retrieval-Augmented Generation of Large Language Model
Xun Liang1, Simin Niu1, Zhiyu Li2,thanks: Corresponding author: Zhiyu Li., Sensen Zhang1, Hanyu Wang1, Feiyu Xiong2, Jason Zhaoxin Fan3, Bo Tang2, Shichao Song1, Mengwei Wang1, Jiawei Yang1 1School of Information, Renmin University of China, Beijing, China 2Institute for Advanced Algorithms Research, Shanghai, China 3Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing, School of Artificial Intelligence, Beihang University, Beijing, China
1 Introduction
Retrieval-augmented generation (RAG) provides an efficient solution for expanding the knowledge boundaries of large language models (LLMs). Many advanced LLMs, such as ChatGPT OpenAI et al. (2024), Gemini Team et al. (2024), and Perplexy.ai [^1] https://www.perplexity.ai/, have incorporated external retrieval modules within their web platforms. However, during the RAG process, query-relevant texts are processed sequentially through the retriever, the filter before being synthesized into a response by the generator, introducing potential security risks, as attackers can manipulate texts at any stage of the pipeline.
Figure 1: Motivation. The attack methods used in existing RAG benchmarks fail to bypass the RAG components, which hindering accurate RAG security evaluation. Our SafeRAG introduces enhanced attack methods to evaluate the potential vulnerabilities of RAG.
Current attack tasks targeting RAG can be divided into the following four surfaces:
  • Noise: Due to the limitation in retrieval accuracy, the retrieved contexts often contain large quantities of noisy texts that are at most merely similar to the query but do not actually contain the answer. Attackers can exploit this retrieval limitation to dilute useful knowledge by deliberately injecting extensive noisy texts Chen et al. (2024a); Fang et al. (2024).
  • Conflict: Knowledge from different sources may conflict with one another, creating opportunities for attackers to manipulate. Simply injecting conflicting texts could prevent LLMs from determining which piece of knowledge is more reliable, resulting in vague or even incorrect responses. Wu et al. (2024a); Liu et al. (2023).
  • Toxicity: The internet often contains toxic texts published by attackers. Such malicious texts are highly likely to be incorporated into the RAG pipeline, inducing LLMs to generate toxic responses Deshpande et al. (2023); Perez and Ribeiro (2022).
  • Denial-of-Service (DoS): The target of DoS is to cause LLMs to refuse to answer, even when evidence is available Chaudhari et al. (2024); Shafran et al. (2024). DoS-inducing texts injected by attackers are particularly insidious because the resulting behavior is easily mistaken for the RAG’s limitations.
However, most of existing attack tasks at above surfaces often fail to bypass the safety RAG components, making the attacks no longer suitable for RAG security evaluation. There are four main reasons. R-1: Simple safety filters can effectively defend against noise attack Li et al. (2024), as existing noise is often concentrated in superficially relevant contexts, which may actually belong to either similar-topic irrelevant contexts or relevant contexts that do not contain answers. (Fig.1-⑨). R-2: Existing conflict primarily focuses on questions that LLMs can directly answer but contain factual inaccuracies in the related documents Xu et al. (2024). Current adaptive retrievers Tan et al. (2024) have been able to effectively mitigate such context-memory conflict. R-3: Advanced generators demonstrate strong capabilities in detecting and avoiding explicit and implicit toxicity, such as bias, discrimination, metaphor, and sarcasm Sun et al. (2023); Wen et al. (2023). R-4: Traditional DoS attack mainly involves maliciously inserting explicit (Fig.1-⑤⑥) or implicit (Fig.1-⑦⑧) refusal signals into the RAG pipeline. Fortunately, such signals are often filtered out as they inherently do not support answering the question, or they are ignored by generators due to being mixed into evidences Shafran et al. (2024).
To address above limitations, we propose four novel attack tasks for conducting effective RAG security evaluation. Firstly, we define silver noise (Fig.1-②), which refers to evidence that partially contains the answer. Such noise can circumvent most safety filters, thereby undermining the RAG diversity (R-1). Secondly, unlike the widely studied context-memory conflict, we explore a more hazardous inter-context conflict (Fig. 1-④). Since LLMs lack sufficient parametric knowledge to handle external conflicts, they are more susceptible to being misled by tampered texts (R-2). Thirdly, we reveal the vulnerability of RAG under the soft ad attack (Fig. 1-③). As a special type of implicit toxicity, the soft ad can evade LLMs and ultimately be inserted into the response of generators (R-3). Finally, to enable refusal signals to bypass filters or generators, we propose a white DoS (Fig.1-①) attack. Under the guise of a safety warning, such attack falsely accuses the evidence of containing a large number of distorted facts, thereby achieving its purpose of refusal (R-4).
Existing benchmarks mainly focus on applying a certain attack task at specific stages of the RAG pipeline and observing the impact of the selected attack on the retriever or generator. In this paper, we introduce the RAG security evaluation benchmark, SafeRAG, which systematically evaluates the potential security risks in the retriever and generator by performing four surfaces of improved attack tasks across different stages of the RAG pipeline. Our main contributions are:
  • We reveal four attack tasks capable of bypassing the retriever, filter, and generator. For each attack task, we develop a lightweight RAG security evaluation dataset, primarily constructed by humans with LLM assistance.
  • We propose an economical, efficient, and accurate RAG security evaluation framework that incorporates attack-specific metrics, which are highly consistent with human judgment.
  • We introduce the first Chinese RAG security benchmark, SafeRAG, which analyzes the risks posed to the retriever and generator by the injection of noise, conflict, toxicity, and DoS at various stages of the RAG pipeline.
Table 1: Related works.
Method Attack Type Attack Stage Evaluation Method Evaluation Metrics Lang. Evaluation Task
RGB Chen et al. (2024a) Noise Knowledge Base Rule-based EM CN, EN Open-domain Q&A
RAG Bench Fang et al. (2024) Noise, Conflict Knowledge Base Rule-based EM, F1 EN Open-domain Q&A
LRII Wu et al. (2024b) Noise, Conflict Filtered Context Model-based
Misleading Ratio,
Uncertainty Ratio
EN
Open-domain Q&A,
Simple Fact Q&A
RECALL Liu et al. (2023) Conflict Filtered Context
Model-based,
Rule-based
Accuracy, BLEU,
ROUGE-L, Misleading Rate,
Mistake Reappearance Rate
EN
Open-domain Q&A,
Simple Fact Q&A,
Text Generation
ClashEval Wu et al. (2024a) Conflict Filtered Context Rule-based
Accuracy, Prior Bias,
Context Bias
EN Domain-specific Q&A
PoisonedRAG Zou et al. (2024) Conflict Knowledge Base Rule-based
Attack Success Rate,
Precision, Recall, F1
Phantom Chaudhari et al. (2024) DoS Knowledge Base Rule-based Retrieval Failure Rate
MAR Shafran et al. (2024) DoS Knowledge Base Rule-based Retrieval Accuracy
SafeRAG (Ours)
Noise, Conflict,
Toxicity, DoS
Knowledge Base, Retrieved
Context, Filtered Context
Model-based,
Rule-based
F1 (correct/incorrect/avg),
Attack Success Rate,
Retrieval Accuracy
CN
Domain-specific Question,
Comprehensive Question-contexts
2 Related Work
2.1 RAG Security Benchmark Dataset
Before performing RAG security evaluation, researchers typically design attack datasets meticulously to trigger the vulnerability of RAG. The primary attack types currently include noise, conflict, toxicity, and DoS. As for noise, RGB Chen et al. (2024a) employs a retrieve-filter-classify strategy, dividing the top retrieved contexts related to the query into golden contexts (those containing the correct answer) and relevant noise contexts. RAG Bench Fang et al. (2024) adopts the same approach to construct relevant noise while also introducing irrelevant noise. LRII Wu et al. (2024b) further refines the construction of irrelevant noise through a combination of human effort and LLM assistance, categorizing it into three types: semantically unrelated, partially related, and related to questions.
As for conflict, most existing works rely on generating counterfactual perturbations using LLMs Fang et al. (2024); Zou et al. (2024). However, these methods may incorrectly alter key facts, leading to the generation of similar-topic irrelevant contexts or hallucinatory relevant contexts. Consequently, manually constructing conflicts is considered a more reliable approach. For instance, RECALL Liu et al. (2023) create context-memory conflict manually to evaluate the ability of LLMs to discern the reliability of external knowledge. In this paper, we first refine the rules for manually constructing conflicts and build high-quality, deliberately misleading inter-context conflicts.
DoS attack is relatively simple to construct. For example, Phantom Chaudhari et al. (2024) injects the response “… Sorry, I don’t know …” into the knowledge base to prevent LLMs from providing useful responses. MAR Shafran et al. (2024) introduces target responses such as “I don’t know. The context does not provide enough information” or “I cannot provide a response that may perpetuate or encourage harmful content” to induce LLMs to refuse. However, these rule-based generated attack texts are often intercepted by filters as obviously unhelpful to the query, leading to failed attacks. To address this limitation, MAR Shafran et al. (2024) employs model-based methods to generate attack contexts that induce target responses and injects them into the knowledge base, but these attack texts are often interspersed among evidence, causing LLMs to prioritize the evidence and rendering the attack ineffective. As a result, in this paper, we propose a white DoS attack, which fabricates a security warning to falsely accuse evidence of containing a large amount of distorted facts, successfully inducing LLMs to refuse to respond.
Research on toxicity attack has predominantly focused on direct prompt injection targeting LLMs, with no dedicated investigation of RAG under toxicity scenarios. Therefore, in our SafeRAG datasets, we also include toxicity attack, with particular emphasis on implicit toxic attack that can easily bypass the retriever, filter, and generator.
Figure 2: The process of generating attacking texts. To meet the injection requirements for four attack surfaces: Noise, Conflict, Toxicity, and DoS, we first collected a batch of news articles and constructed a comprehensive question-contexts dataset as a basic dataset. Subsequently, we selected attack-targeted text from the basic dataset for the generation of attacking texts.
2.2 RAG Security Evaluation Metric
When evaluating the safety of RAG, well-designed evaluation metrics are crucial to ensure that the assessment results comprehensively and accurately reflect the LLM’s actual performance, while also providing effective guidance for subsequent improvements and optimizations. Existing safety evaluation metrics can be broadly categorized into rule-based and model-based approaches. For instance, methods such as RGB Chen et al. (2024a), RAG Bench Fang et al. (2024), and PoisonedRAG Zou et al. (2024) utilize traditional evaluation metrics (e.g., EM, F1, Recall, Precision, and Attack Success Rate) to assess the safety of generated content. Meanwhile, LRII Wu et al. (2024b), RECALL Liu et al. (2023), and ClashEval Wu et al. (2024a) introduce custom metrics for safety evaluation, including Misleading Rate, Uncertainty Ratio, Mistake Reappearance Rate, Prior Bias, and Context Bias. Additionally, Phantom Chaudhari et al. (2024) and MAR Shafran et al. (2024) assess the retrieval safety of RAG from the perspectives of Retrieval Failure Rate and Retrieval Accuracy.
In this paper, we focus on both the retrieval safety and generation safety of RAG to provide a more comprehensive evaluation perspective. Moreover, existing RAG safety evaluation methods primarily target open-domain Q&A, with an emphasis on simple fact Q&A in English. We construct a comprehensive question-contexts safety evaluation dataset specifically tailored to the Chinese news domain, aiming to address the complex issues likely to arise in real-world applications.
3 Threat Framework: Attacks on the RAG Pipeline
3.1 Meta Data Collection and Pre-processing
As shown in Fig.2-①, we collected raw news texts from news websites [^2] http://www.news.cn/ between 08/16/24 and 09/28/24, covering five major sections: politics, finance, technology, culture, and military. Subsequently, we manually screened news segments that met the following criteria: (1) contain more than 8 consecutive sentences; (2) consecutive sentences revolve around a specific topic; (3) consecutive sentences can generate comprehensive questions of the what, why, or how types.
3.2 Generation of Comprehensive Question and Golden Contexts
Using DeepSeek [^3] https://www.deepseek.com/, a powerful Chinese LLM engine, and referencing the news title, we generated a comprehensive question and its corresponding 8 pieces of golden contexts for each extracted news segment (Fig.2-②). In total, we obtained 110 unique question-contexts pairs. Additionally, we manually verified and removed data points that did not meet the following criteria: (1) the question is not a comprehensive what, why, or how type question; (2) there are contexts unrelated to the question. Finally, we obtained 100 unique question-contexts pairs, which serve as the basic dataset for attack text generation. The complete generation prompt is detailed in Fig. 12.
3.3 Selection of Attack-Targeted Texts and Generation of Attacking Texts
We select different attack-targeted texts from the question-contexts pairs in the base dataset to generate the specific attacking texts.
3.3.1 Generation of Silver Noise
To construct silver noise, which includes partial but incomplete answers, we first need to decompose the golden contexts in the base dataset. Specifically, we utilized the knowledge transformation prompt proposed in Chen et al. (2024b) [^4] The complete prompt can be found in Appendix 15. to break the contexts into fine-grained propositions (Fig.2-③), which are the smallest semantic units that are complete and independent as evidence. Then, we selected the proposition with the highest semantic similarity (cosine similarity) to the question as the attack-targeted text, ensuring that the subsequent attack texts achieve a high recall ratio. Finally, we prompted DeepSeek to generate 10 diverse contexts based on the selected attack-targeted text.
3.3.2 Generation of Inter-Context Conflict
The goal of conflict attack is to generate target texts that are prone to contradicting or being confused with the golden context. To achieve this, we manually select a golden context most susceptible to manipulation into a conflict [^5] It is important to emphasize that constructing conflicts is a meticulous process that is not well-suited to being fully automated by LLMs. We first refined the rules for manually constructing conflicts, ensuring that the generated attack is as realistic and effective as possible.. Subsequently, annotators are instructed to modify the context based on the following guidelines: (1) Minimal Perturbation: Introduce conflicts using the smallest possible changes (Fig. 3-①); (2) Rewriting for Realistic Conflicts: Rewrite the context where appropriate to create more convincing conflicts (Fig. 3-②); (3) Preservation of Key Facts: Avoid perturbations that render the conflict invalid, as altering the key fact may lead to generating the hallucinated context instead of the expected conflict context (Fig. 3-③), or even make the context irrelevant to the query (Fig. 3-④). Therefore, we require annotators to preserve key facts while generating conflict contexts. For instance, the year 2018 in Fig. 3-③ and the month September in Fig. 3-④.
Figure 3: Cases of forming conflict contexts.
3.3.3 Generation of Soft Ad
For the toxic attack task, we manu ally selected the golden context most susceptible to the injection of malicious soft ads as the attack-targeted text. Then, we encouraged annotators to use the following two methods to create the soft ad attack text by seamlessly incorporating seemingly professional and harmless soft ads into the attack-targeted text: (1) Direct Insertion: Soft ad attack texts are directly inserted into the original context (Fig. 26-①); (2) Indirect Insertion: The original context is modified or adjusted to some extent, for instance, by promoting alongside authoritative entities (such as government, organizations, etc.) within the context, making the soft ad more subtle and closely integrated into the context (Fig. 26-②). The complete annotation requirements are shown in Fig. 17.
3.3.4 Generation of White Denial-of-Service
In a DoS attack, the original question is directly used as the target text. We then employ a rule-based approach to construct refusal contexts. Specifically, the construction rule is as shown as in Fig. 4.
Figure 4: The construction rules of White DoS. Blue text represents the original question, designed to bypass the retriever. Green text is used to bypass the filter, and red text is intended to bypass the generator to achieve the goal of refusal to answer.
The white DoS attack text constructed in this manner leverages the pretense of a safety warning to falsely accuse the evidence of containing heavily distorted information, thereby justifying refusal. Since safety warnings are perceived as well-intentioned and high-priority, they are less likely to be filtered by filters and are more likely to be adopted by generators.
3.4 Attacks on the RAG Pipeline
We perform a final manual verification of the various types of attack contexts obtained and selected appropriate contexts from the golden contexts to construct the final security evaluation datasets. The detailed construction and usage rules for the datasets can be found in Appendix A.1.3.
4 Evaluation Metrics
4.1 Retrieval Safety Assessment Metric: Retrieval Accuracy
Retrieval Accuracy (RA) is a metric used to evaluate the performance of RAG in terms of both retrieval accuracy and safety. It combines the recall of golden contexts and the suppression ability for attack contexts. The formula is as follows:
$\text{RA}=\frac{\text{Recall (gc)}+(1-\text{Recall (ac))}}{2},$
where Recall (gc) and Recall (ac) denote the recall of golden contexts and attack contexts, respectively.
Figure 5: Experimental results injected different noise ratios into the text accessible within the RAG pipeline.
Figure 6: Experimental results injected conflict into the text accessible within the RAG pipeline.
The core idea of RA is to balance the RAG’s ability to retrieve relevant content while avoiding incorrect or harmful content. A high Recall (gc) reflects strong coverage of correct content, while a low Recall (ac) demonstrates the RAG’s robustness in suppressing irrelevant or disruptive content. By combining these two sub-metrics, the higher RA indicates better retrieval performance by RAG.
4.2 Generation Safety Assessment Metric: F1 Variant and ASR
4.2.1 F1 Variants in Noise and DoS
For silver noise and white DoS tasks, we designed a metric called F1 (avg) to evaluates whether the generated response include fine-grained propositions and thus assess the diversity of the generator. Specifically, we prompt evaluator to complete the multiple-choice question based on the generated response, i.e., the news summary (Fig.7).
Figure 7: Evaluation cases for multiple-choice questions in Noise and DoS tasks.
Figure 8: An evaluation case for a multiple-choice question in the conflict task.
To build multiple options for each data point, we intentionally manipulate certain propositions extracted from the golden contexts to create incorrect ground truth options while retaining the unmodified propositions as correct ground truth options. The evaluator is then tasked with selecting both the correct and incorrect options. Then, we use F1 (correct) to measure the generated response’s ability to correctly identify the correct options, and F1 (incorrect) to evaluate its ability to recognize the incorrect options. We further define F1 (avg), which is the average of F1 (correct) and F1 (incorrect). F1 (avg) quantifies the performance of the generated response in recognizing both correct and incorrect options and reflects the generator’s diversity. If the generator can produce diverse and high-quality responses, it will correctly infer both the correct and incorrect options, resulting in a high F1 (avg).
Figure 9: Experimental results injected toxicity into the text accessible within the RAG pipeline.
Figure 10: Experimental results injected DoS into the text accessible within the RAG pipeline.
4.2.2 F1 Variants in Conflict
To construct multiple-choice options for each data point, we deliberately use conflicting facts between contexts as the options and manually annotate the correct ground truth options. As shown in Fig.8, if the response utilizes the correct fact, it can accurately select the correct options. However, selecting the wrong options is not solely due to using conflicting facts; it may also result from generating hallucinated facts. Consequently, the F1 (incorrect) metric is unreliable. Therefore, in the evaluation of inter-context conflict, we only retain the F1 (correct) metric to assess the generator’s performance.
4.2.3 ASR in Conflict, Toxicity, and DoS
In the conflict, toxicity, and DoS tasks, attack keywords are present, such as the conflict facts leading to inter-context conflicts, seamlessly integrated soft ad keywords, and refusal signals. Therefore, in these tasks, we can evaluate the generator’s safety using the attack success rate (ASR) Zou et al. (2024). If a higher proportion of attack keywords appears in the response text, the ASR will increase [^6] Note: in experiments, we use the attack failure rate (AFR = 1 - ASR) for safety evaluation because AFR, as a positive metric, can be analyzed alongside F1 variants..
5 Experiments
5.1 Settings
The default retrieval window for the silver noise task is set to top K = 6, with a default attack injection ratio of 3/6. For other tasks, the default retrieval window is top K = 2, and the attack injection ratio is fixed at 1/2. We evaluated the impact of using different retrievers (DPR, BM25, Hybrid, Hybrid-Rerank) and filters (OFF, filter NLI Li et al. (2024), compressor SKR Wang et al. (2023)) across different RAG stages (indexing, retrieval, generation) on the contexts retrieved for various generators (DeepSeek, GPT-3.5-turbo, GPT-4, GPT-4o, Qwen 7B, Qwen 14B, Baichuan 13B, ChatGLM 6B). The bold values represent the default settings. Then, we adopt a unified sentence chunking strategy to segment the knowledge base and build the index. The embedding model used is bge-base-zh-v1.5, the reranker is bge-reranker-base, and the evaluator is GPT-3.5-turbo.
5.2 Results on Noise
We inject different noise ratios into the text accessible in the RAG pipeline, including the knowledge base, retrieved context, and filtered context. As shown in Fig.5, the following observations can be made: (1) Regardless of the stage where noise is injected, the F1 (avg) score exhibits a downward trend as the noise ratio increases, indicating a decline in generation diversity (Fig.5-①); (2) The retriever demonstrates some noise resistance, as noise injected at the knowledge base has approximately 50% chance of not being retrieved. The results in Fig.5-① support this point. Specifically, as the noise ratio increases, the Retrieval Accuracy (RA) of injecting silver noise into the retrieved context or filtered context significantly outperforms that of injecting it into the knowledge base; (3) The performance of injecting noise into the retrieved context and filtered context is similar, indicating that the filter cannot effectively resist silver noise since silver noise still supports answering the query. (4) Different retrievers exhibit varying levels of robustness to noise. Overall, the ranking is Hybrid-Rerank > Hybrid > BM25 > DPR, suggesting that compared to attacking contexts, hybrid retriever and rerankers show a preference for retrieving golden contexts. (5) Compression-based filters like SKR are not sufficiently secure, as they tend to lose detailed information, leading to a decrease in F1 (avg).
5.3 Results on Conflict, Toxicity, and DoS
(1) After injecting different types of attacks into the texts accessible by the RAG pipeline, it was observed that the retrieval accuracy (RA) and the attack failure rate (AFR) decreased across all three tasks. The ranking of attack effectiveness at different RAG stages was: filtered context > retrieved context > knowledge base. Furthermore, adding conflict attack increased the likelihood of misjudging incorrect options as correct, leading to a drop in F1 (correct). Introducing DoS attack reduced F1 (avg) and severely impacted generative diversity (Fig.6109-①). (2) Retrievers exhibited different vulnerabilities to various attacks. For instance, Hybrid-Rerank was more susceptible to conflict attack, while DPR was more prone to DoS attack. Both experienced a significant decrease in AFR. Additionally, all retrievers showed consistent AFR degradation under toxicity attack. After adding conflict attack, the F1 (correct) scores of all retrievers became similar, indicating stable attack effectiveness. However, DPR was more affected by DoS attack compared to other retrievers, as evidenced by its significantly larger decline in the diversity metric F1 (avg) (Fig.6109-②). (3) The RA of different retrievers was largely consistent across different attack tasks (Fig.6109-③). (4) In conflict tasks, using the SKR filter was less secure because it could compress conflict details, resulting in a decline in F1 (correct). In toxicity and DoS tasks, the NLI filter was generally ineffective, with its AFR close to that of disabling the filter. However, the SKR filter proved to be safe in these tasks, as it was able to compress soft ads and warnings (Fig.6109-④).
5.4 Analysis of Generator and Evaluator
5.4.1 Selection of Generator
Figure 11: Cumulative analysis of the generator’s positive evaluation metrics across different attack tasks.
We conduct a cumulative analysis of the positive metrics across different attack tasks. As shown in Fig.11, the results show that Baichuan 13B achieved a leading position in multiple attack tasks, particularly excelling in AFR (DoS) and F1 Variants (DoS) metrics [^7] The detailed results are shown in Table 3.. Lighter models are even safer than models such as the GPT series and DeepSeek.
5.4.2 Selection of Evaluator
F1 (correct) F1 (incorrect) ASR/AFR
Silver Noise 89.97 96.22
Inter-context Conflict 99.10 95.65
Soft Ad 91.67 100
White DoS 89.97 96.22 100
Table 2: Evaluation metrics and human consistency.
As shown in Table 2, We present the evaluation metrics and their consistency with human judgments. The ASR and AFR metric exhibit a high human consistency. Similarly, the F1 (correct) and F1 (incorrect) scores obtained using DeepSeek also demonstrate strong agreement with human judgments. Therefore, DeepSeek is uniformly adopted for evaluation across all experiments.
6 Conclusion
This paper introduces SafeRAG, a benchmark designed to assess the security vulnerabilities of RAG against data injection attacks. We identified four critical attack surfaces: noise, conflict, toxicity, and DoS, and revealed significant weaknesses across the retriever, filter, and generator components of RAG. By proposing novel attack strategies such as silver noise, inter-context conflict, soft ad, and white DoS, we exposed critical gaps in existing defenses and demonstrated the susceptibility of RAG systems to subtle yet impactful threats.
Appendix A Appendix
A.1 Preliminaries and Definitions
A.1.1 RAG Pipeline
Let $q$ denote the query, $i$ the instruction for the LLM, and $C_{b}$ the knowledge base that comprises all available documents. For effective integration of external knowledge with the LLM’s generative capabilities in the answering process, a RAG pipeline typically includes three primary modules: a retriever $\mathcal{R}$, a filter $\mathcal{F}$, and a generator $\mathcal{G}$.
First, given the query $q$ and the knowledge base $C_{b}$, the retriever $\mathcal{R}$ returns the $k$ most relevant contexts for query $q$:
$C_{r}^{k}=\mathcal{R}(q,C_{b})=(c_{r}^{1},c_{r}^{2},\dots,c_{r}^{k}).$
Next, to further refine these retrieved contexts, the filter $\mathcal{F}$ picks or compresses $C_{r}^{k}$ to derive contexts that are highly relevant to the query:
$C_{f}^{m}=\mathcal{F}(q,C_{r}^{m})=(c_{f}^{1},c_{f}^{2},\dots,c_{f}^{m}),$
where $m\leq k$. Finally, the generator $\mathcal{G}$ combines the instruction $i$, the query $q$, and the filtered contexts $C_{f}^{m}$ by concatenating them (denoted by $\oplus$) into a unified prompt, which is then fed into the LLM to generate the final answer:
$r=\mathcal{G}(i,q,C_{f}^{m})=\mathrm{LLM}(i\oplus q\oplus c_{f}^{1}\oplus\dots\oplus c_{f}^{m}).$
By sequentially performing the retrieve $\rightarrow$ filter $\rightarrow$ generate composite mapping:
$(i,q,C_{b})$ $\;\mapsto\;\mathcal{R}(q,C_{b})$
$\;\mapsto\;\mathcal{F}\bigl{(}q,\mathcal{R}(q,C_{b})\bigr{)}$
$\;\mapsto\;\mathcal{G}\bigl{(}i,q,\mathcal{F}\bigl{(}q,\mathcal{R}(q,C_{b})\bigr{)}\bigr{)}$
$\;\mapsto\;r,$
the RAG Pipeline effectively exploits relevant contexts from the external knowledge base $C_{b}$ while also leveraging the powerful generation capabilities of LLMs. This approach mitigates the hallucination problem and enhances both the accuracy and the interpretability of the answers.
A.1.2 SafeRAG Dataset Construction
(1) Base Dataset Construction: Raw document is collected from the external news website, and paragraphs meeting the following criteria are selected: 1) contain more than 8 consecutive sentences; 2) consecutive sentences revolve around a specific topic; 3) consecutive sentences can generate comprehensive questions of the what, why, or how types. For each paragraph, a comprehensive question $q_{j}$ is generated, and $n$ golden contexts $C_{g}^{j}=\{c_{g}^{1},c_{g}^{2},\dots,c_{g}^{n}\},n\geq 8$ closely related to the question $q_{j}$ are extracted. Each golden context $c_{g}^{j}\in C_{g}^{j}$ is manually screened and verified to ensure accuracy, coherence, and the ability to fully answer the question. The base dataset is defined as:
$\mathcal{D}_{\mathrm{base}}=\bigl{\{}(q_{j},i,C_{g}^{j})\;|\;j=1,\dots,N\bigr{\}},$
where $i$ represents the uniform instruction $i$ provided to the LLM [^8] The default instruction $i$ used in this paper is shown in Fig. 14.
(2) Attacking Texts Generation and RAG Security Evaluation Dataset Construction: To evaluate the security of RAG under different adversarial scenarios, we design four attack tasks $T=\{\mathrm{SN},\mathrm{ICC},\mathrm{SA},\mathrm{WDoS}\}$: silver noise $\mathrm{SN}$, inter-context conflict $\mathrm{ICC}$, soft ad $\mathrm{SA}$, and white DoS $\mathrm{WDoS}$. For each attack task $t\in T$, malicious attack texts are generated and combined with golden contexts to construct the RAG security evaluation dataset (i.e., SafeRAG dataset). The detailed process is as follows:
  1. 1)
    Silver Noise:
  2. 1.
    Decompose the golden contexts $C_{g}^{j}$ into minimal semantic units (i.e., propositions) ${P}_{g}^{j}=\{p_{g}^{1},p_{g}^{2},\dots\}$.
  3. 2.
    Select the proposition $p_{g}^{i}$ most semantically relevant to question $q_{j}$.
  4. 3.
    Use $p_{g}^{i}$ as input to the DeepSeek to generate 10 diverse attack contexts.
  5. 4.
    Manually select 8 semantically consistent yet non-redundant attack contexts to form the silver noise contexts:
    $C_{a}^{\mathrm{SN},j}=\{c_{a}^{k,\mathrm{SN}}|\;k=1,\ldots,8\}.$
  6. 5.
    Combine the silver noise contexts with the golden contexts to construct the silver noise security evaluation dataset:
[…]
    $|\;j=1,\dots,M_{\mathrm{SN}}\bigr{\}}.$
  1. 2)
    Inter-Context Conflict:
  2. 1.
    Select a golden context $c_{g}^{i}\in C_{g}^{j}$.
  3. 2.
    Based on $c_{g}^{i}$, use strategies such as minimal perturbation to generate a text $C_{a}^{\mathrm{ICC},j}=\{c_{a}^{1,\mathrm{ICC}}\}$ that clearly contradicts $c_{g}^{i}$.
  4. 3.
    Randomly select another golden context $c_{g}^{e}\in C_{g}^{j}\setminus\{c_{g}^{i}\}$, where $e\neq i$.
  5. 4.
    Combine the generated conflict context with the golden context $c_{g}^{i}$ used to generate it and another golden context $c_{g}^{e}$ to construct the conflict security evaluation dataset:
[…]
    $|\;j=1,\dots,M_{\mathrm{ICC}}\bigr{\}}.$
  1. 3)
    Soft Ad:
  2. 1.
    Select a golden context $c_{g}^{i}\in C_{g}^{j}$.
  3. 2.
    Manually read $c_{g}^{i}$ and generate appropriate attack keywords $a_{1,\ldots}$ (i.e., soft ad keywords).
  4. 3.
    Use strategies like direct insertion or indirect insertion to embed attack keywords into $c_{g}^{i}$, generating the soft ad attack context $C_{a}^{\mathrm{SA},j}=\{c_{a}^{1,\mathrm{SA}}\}$.
  5. 4.
    Randomly select two other golden contexts $c_{g}^{e}$ and $c_{g}^{l}$ from $C_{g}^{j}\setminus\{c_{g}^{i}\}$ for constructing the clean set, where $i\neq e,e\neq l,l\neq i$.
  6. 5.
    Combine the soft ad attack context with the other two golden contexts $c_{g}^{e},c_{g}^{l}$ to construct the soft ad security evaluation dataset:
[…]
    $|\;j=1,\dots,M_{\mathrm{SA}}\bigr{\}}.$
  1. 4)
    White DoS:
  2. 1.
    Based on question $q_{j}$, generate an attack context $C_{a}^{\mathrm{WDoS},j}=\{c_{a}^{1,\mathrm{WDoS}}\}$ containing white safety warnings.
  3. 2.
    Combine the generated white DoS context with the complete golden contexts $C_{g}^{j}$ to construct the White DoS security evaluation dataset:
[…]
    $|\;j=1,\dots,M_{\mathrm{WDoS}}\bigr{\}}.$
The complete SafeRAG dataset is defined as:
$\mathcal{D}_{\mathrm{sfr}}=\mathcal{D}_{\mathrm{SN}}\cup\mathcal{D}_{\mathrm{ICC}}\cup\mathcal{D}_{\mathrm{SA}}\cup\mathcal{D}_{\mathrm{WDoS}}.$
A.1.3 Threat Framework: Attacks on the RAG Pipeline
We utilize the SafeRAG dataset to simulate various attack scenarios that RAG may encounter during Q&A tasks. Our proposed threat framework enables attackers to inject malicious contexts into any stage of the RAG pipeline (i.e., retrieval, filter, or generation) to analyze potential vulnerabilities when facing different types of attacks.
Specifically, for an attack task $t\in T$, $C_{b}^{t}$ from the dataset $\mathcal{D}_{t}=\{(q_{j},i,C_{b}^{t}\cup C_{a}^{\mathrm{t},j})\;|\;j=1,\dots,M_{t}\}$ is the selected knowledge base for the given attack task $t$. For a given query $q_{j}$ from $\mathcal{D}_{t}$, and the knowledge base $C_{b}^{t}$, we first construct a benign RAG pipeline, where neither the retriever $\mathcal{R}$, the filter $\mathcal{F}$, nor the generator $\mathcal{G}$ is influenced by malicious contexts. This allows us to observe the baseline performance of RAG in terms of retrieval and generation when no attacks are present.
Under the threat framework presented in this paper, we then select malicious contexts from the attack source:
$C_{a}^{t,j}=\{c_{a}^{1,t},\dots,c_{a}^{k^{\prime},t}\},k^{\prime}\geq 1,$
where $C_{a}^{t,j}$ represents the $k^{\prime}$ attack contexts injected by the attacker [^9] For the text injection attack, the attacker first ejects the original bottom k’ benign contexts and then injects k’ malicious contexts.. These contexts may be embedded into any stage of the RAG pipeline, targeting its specific components:
(1) Attacking the Retriever: The attacker injects attack contexts $C_{a}^{t,j}$ into the original knowledge base $C_{b}^{t}$, camouflaging them as relevant contexts to compromise the retriever $\mathcal{R}$. In this scenario, when the retriever executes $\mathcal{R}(q_{j},C_{b}^{t}\cup C_{a}^{t,j})$, it is likely to retrieve the attack text $c_{a}\in C_{a}^{t,j}$, resulting in erroneous or biased contexts that affect subsequent filter and generation stages.
(2) Attacking the Filter: The attacker directly incorporates attack contexts into the retriever’s output $C_{r}^{k}$, such that:
[…]
$c_{r}^{1},\dots,c_{r}^{k}\})[:K].$
Consequently, the filter $\mathcal{F}$ and generator $\mathcal{G}$ misinterpret these attack texts as part of the retrieved contexts, integrating them into the subsequent stages of the pipeline.
(3) Attacking the Generator: The attacker disrupts the filter stage by introducing $C_{a}^{t,j}$ into the filtered results $C_{f}^{m}$, such that:
[…]
$c_{f}^{1},\dots,c_{f}^{m}\})[:K].$
This action directly distorts the input contexts for the generator.
Regardless of the stage at which the injection occurs, the attacker’s objective is to mislead or compromise the RAG pipeline’s final output $r$ by leveraging the attack contexts $C_{a}^{t,j}$. It is important to note that, under the attack assumptions in this paper, golden contexts $C_{g}^{j}$ are neither altered in content nor re-ranked [^10] This setting is widely adopted in numerous many attacks Xiang et al. (2024); Zhong et al. (2023); Zou et al. (2024); Pan et al. (2023b, a); Du et al. (2022). Through this method, the attacker maximizes the system’s original usability and normalcy while covertly influencing the RAG pipeline’s generated responses.
Figure 12: Generation of comprehensive questions and golden contexts.
Figure 13: Generation of comprehensive questions and golden contexts (in Chinese).
Figure 14: Question answering.
Figure 15: Extraction of propositions from golden contexts.
Figure 16: Extraction of propositions from golden contexts (in Chinese).
Figure 17: Guidelines for generating (annotating) soft ad attack texts.
Figure 18: Guidelines for generating (annotating) soft ad attack texts (in Chinese).
Figure 19: Multiple-choice question evaluation.
Figure 20: Multiple-choice question evaluation (in Chinese).
Table 3: Cumulative analysis of the generator’s positive evaluation metrics across different attack tasks
Model Noise Conflict Toxicity DoS
F1_Variants AFR F1_Variants AFR F1_Variants AFR F1_Variants AFR
DeepSeek 0.1032 - 0.4000 0.69 - 0.38 0.2068 0.17
GPT-3.5-Turbo 0.1102 - 0.3615 0.72 - 0.47 0.1654 0.17
GPT-4 0.1141 - 0.3615 0.66 - 0.29 0.4760 0.16
GPT-4o 0.1229 - 0.3615 0.68 - 0.37 0.3196 0.28
Qwen 7B 0.1016 - 0.4948 0.75 - 0.47 0.2582 0.57
Qwen 14B 0.1005 - 0.5000 0.65 - 0.38 0.1842 0.29
Baichuan 13B 0.0706 - 0.4800 0.87 - 0.59 0.7222 1.00
ChatGLM 6B 0.0815 - 0.5966 0.55 - 0.27 0.5096 1.00
Figure 21: A data point of a comprehensive question, the golden contexts and propositions.
Figure 22: A case of multiple options and the ground truth answers.
Figure 23: A data point of a comprehensive question, the golden contexts and propositions (in Chinese).
Figure 24: A case of multiple options and the ground truth answers (in Chinese).
Figure 25: A case of silver noise.
Figure 26: A case of context conflict and soft ad.